OpenStudy (kayders1997):

Will someone help me with this integral problem, evaluate the integral from 1 to 3 when the equation is y^5-2y/y^3 dy

OpenStudy (harman.singh):

Just to be sure, is this what the question looks like? \[\int\limits_{3}^{1}\frac{ y^5-2y }{ y^3 }.dy\]

OpenStudy (kayders1997):

3 is on the top but otherwise yes

OpenStudy (harman.singh):

Alright. So you can integrate it by dividing it into two separate fractions \[\int\limits_{1}^{3}\frac{ y^5 }{ y^3 } - \int\limits_{1}^{3}\frac{ 2y }{ y^3 }\]

OpenStudy (kayders1997):

Alright

OpenStudy (kayders1997):

So you will have integral of y^2-2/y from 3 to 1?

OpenStudy (harman.singh):

Yes, just one thing. It would be 2/y^2 for the second one

OpenStudy (harman.singh):

\[\frac{ 2y }{ y^3 }=\frac{ 2 }{ y^2 }\]

OpenStudy (kayders1997):

Sorry I'm back

OpenStudy (kayders1997):

@harman.singh

OpenStudy (harman.singh):

Thats fine. I am here

OpenStudy (kayders1997):

So than antiderivative of that?

OpenStudy (harman.singh):

Yes, correct

OpenStudy (kayders1997):

So than that would be (1/3)y^3 and lets see you can rewrite the other one to -2y^-2 so is it 2/y?

OpenStudy (kayders1997):

I think I'm not sure about the second part though

OpenStudy (harman.singh):

Yes, correct so far :)

OpenStudy (kayders1997):

Awesome!! :)

OpenStudy (harman.singh):

So now, it looks like \[\frac{ y^3 }{ 3 }+\frac{ 2 }{ y }+C\]

OpenStudy (harman.singh):

All you need to do now is compute the boundaries for 3 & 1 to get your answer

OpenStudy (kayders1997):

Ok I'll let you know what I get!

OpenStudy (harman.singh):

Sure :)

OpenStudy (kayders1997):

Is it 7.33333?

OpenStudy (harman.singh):

That's the right answer. Good job! :)

OpenStudy (kayders1997):

Thank you so much!!!

OpenStudy (harman.singh):

My pleasure!

OpenStudy (kayders1997):

:)